What is the radius of the circle between and tangent to the circles with these equations:
(x+1)^2 + (y-7)^2=1
(x-6)^2 + (y-1)^2=1

Answer by lwsshak3(11628)

What is the radius of the circle between and tangent to the circles with these equations:
(x+1)^2 + (y-7)^2=1
(x-6)^2 + (y-1)^2=1
***
circle A with center at (-1,7)
circle B with center at (6,1)
circle C (in-between circle
Radius of circle C=(1/2)(distance between centers of circles A and B minus the sum of their radii)
use distance formula:d=√(x1-x2)^2+(y1-y2)^2)
=√(-1-6)^2+(7-1)/2)=√(49+36)=√85
Radius of circle C=(1/2)(√85-2)≈3.61